OpenStudy (anonymous):

x + y = 4 x - 2 y = -5 If the first equation is multiplied by 2 and then the equations are added, the result is _____. 3 x = 13 3 x = 3 x = 3

OpenStudy (anonymous):

still a bit confused but hopefully i can finish this next one

OpenStudy (solomonzelman):

subtract the equations from each other.

OpenStudy (shamim):

3x=3 x=3/3 x=?

OpenStudy (anonymous):

still a bit confused

OpenStudy (anonymous):

dude man

OpenStudy (shamim):

8-5=?

OpenStudy (solomonzelman):

wrergvsvsavbbre

OpenStudy (anonymous):

i can see why because it is a ridiculous questions they should just ask you to solve it, not ask about the steps along the way

OpenStudy (anonymous):

If the first equation is multiplied by 2 and then the equations are added, the result is _____.

OpenStudy (shamim):

3/3=1?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

first step is to multiply the equation by \(2\) all the way across you get \[2x+2y=8\]

OpenStudy (xanthe):